# Unveiling Topological Structures in Text: A Comprehensive Survey of Topological Data Analysis Applications in NLP

### v1

@@ Abstract
The surge of data available on the internet has led to the adoption of various computational methods to analyze and extract valuable insights from this wealth of information. Among these, the field of Machine Learning (ML) has thrived by leveraging data to extract meaningful insights. However, ML techniques face notable challenges when dealing with real-world data, often due to issues of imbalance, noise, insufficient labeling, and high dimensionality. To address these limitations, some researchers advocate for the adoption of Topological Data Analysis (TDA), a statistical approach that discerningly captures the intrinsic shape of data despite noise. Despite its potential, TDA has not gained as much traction within the Natural Language Processing (NLP) domain compared to structurally distinct areas like computer vision. Nevertheless, a dedicated community of researchers has been exploring the application of TDA in NLP, yielding 85 papers we comprehensively survey in this paper. Our findings categorize these efforts into theoretical and non-theoretical approaches. Theoretical approaches aim to explain linguistic phenomena from a topological viewpoint, while non-theoretical approaches merge TDA with ML features, utilizing diverse numerical representation techniques. We conclude by exploring the challenges and unresolved questions that persist in this niche field. Resources and a list of papers on this topic can be found at: https://github.com/AdaUchendu/AwesomeTDA4NLP.  [^1] DISTRIBUTION STATEMENT A. Approved for public release. Distribution is unlimited. This material is based upon work supported by the Department of the Air Force under Air Force Contract No. FA8702-15-D-0001. Any opinions, findings, conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the Department of the Air Force. © 2024 Massachusetts Institute of Technology. Delivered to the U.S. Government with Unlimited Rights, as defined in DFARS Part 252.227-7013 or 7014 (Feb 2014). Notwithstanding any copyright notice, U.S. Government rights in this work are defined by DFARS 252.227-7013 or DFARS 252.227-7014 as detailed above. Use of this work other than as specifically authorized by the U.S. Government may violate any copyrights that exist in this work.
@@ 1 Introduction
TDA aims to answer the main question - what is the true shape of a data? We survey 85 papers that have attempted to find an answer through various approaches. We categorize all observed approaches that incorporate TDA in NLP into two categories, namely theoretical Karlgren et al. (2014); Port et al. (2018) and non-theoretical Zhu (2013); Doshi and Zadrozny (2018). Theoretical approaches involve using TDA to explain linguistic phenomena by probing the topological space, shape, and evolution of topics. On the other hand, Non-theoretical approaches mainly discuss how to effectively apply existing numerical representation techniques in NLP to extract novel topological features with TDA. We will first discuss the principles behind TDA and the two main techniques employed for TDA feature extraction: Persistent Homology and Mapper.
@@ 2.2 Mapper
The intrinsic nature of the Mapper algorithm makes it advantageous in preserving structure even with mapping from one dimension to another. Furthermore, the clustering techniques allow it to be used to explain model performance as the clusters, and colors have meaning that can be further explored. In this survey, we will discuss how several researchers used it to explain several phenomena in NLP tasks.
Finally, Mapper is more useful for exploratory data analysis, while PH is more useful for analyzing point clouds and examining the persistence of features.
@@ 4.1 Tasks, Learning Types, Modality, and TDA techniques
Recently TDA has been more heavily applied to five sub-tasks - deepfake text detection Løvlie (2023); Tulchinskii et al. (2024); Kushnareva et al. (2024, 2021); Uchendu et al. (2024), linguistic/grammatical acceptability classification Cherniavskii et al. (2022); Proskurina et al. (2023); Jain et al. (2024); Perez and Reinauer (2022), word sense induction and disambiguation Rawson et al. (2022); Jakubowski et al. (2020); Tfemčinas (2018), and polysemy word classification Jakubowski et al. (2020); Triki (2021); Shehu (2024); Garcia (2022). All of these five subtasks are particularly interesting because the unique labels have very similar features, making them non-trivial to distinguish from each other. Therefore, researchers employed TDA as it is beneficial for extracting the true structure of a dataset, which could be useful for accurate classification. Lastly, we find that TDA is successful in such applications, achieving comparable performance to the baseline or mostly outperforming baselines.
@@ 4.2.2 Pre-trained Non-contextual Embeddings
Word2Vec embeddings are a type of word representation that allows words with similar meanings to have similar vector representations Mikolov et al. (2013). Word2vec representations have been used as inputs to persistent homology to extract a hierarchical document structure, known as story trees, that allows tracing storylines Haghighatkhah et al. (2022). Word2vec has also been used with mapper to construct a graph of word2vec embeddings to create a topological search engine Cornell (2020). Next, Paluzo Hidalgo et al. (2019) employs TDA with word2vec to measure the distance between the literary style of Spanish poets belonging to similar genres - Francisco de Quevedo, Luis de Góngora, and Lope de Vega. TDA has also been applied on word2vec to solve innovative problems such as the detection of fraudulent papers Tymochko et al. (2021), detecting topological loops in logical statements Tymochko et al. (2020), and disease prediction from epidemic curves Petri and Leitao (2020).
In addition, Wu et al. (2022) using the TDA features extracted from Word2Vec representations, built an ensemble of BERT & TDA to detect contradiction types - Negation, Antonym, Replacement, Switch, Scope, and Latent. Yessenbayev and Kozhirbayev (2022) and Yessenbayev and Kozhirbayev (2024) compare embeddings extracted from text and speech using persistent homology. Holmes (2020), uses TDA to group embeddings for document clustering and topic modeling tasks. Rawson et al. (2022) and Tfemčinas (2018), build an unsupervised topological algorithm for word sense induction and disambiguation. Lastly, Feng et al. (2024) compares the data generated by traditional data augmentation techniques - Word2Vec and GloVe embeddings with LLM-generated augmented data (from GPT-J) using persistent homology.
@@
FastText embeddings build upon the Word2Vec approach by incorporating subword information, enhancing the representation of rare words, and allowing for embedding out-of-vocabulary words Bojanowski et al. (2017). Jakubowski et al. (2020), Triki (2021), and Shehu (2024) use persistent homology to measure the polysemy of a word. They find that topological polysemy is inconsistent across different word embeddings. Jakubowski et al. (2020) further uses a similar technique to perform word sense induction & disambiguation from SemEval-2010 task. Also, Tfemčinas (2018) performs unsupervised word sense induction & disambiguation. Tymochko et al. (2021) extracts topological features from FastText embeddings to detect fraudulent papers.
@@ 4.2.3 Pre-trained Contextual Embeddings
CLS Embedding Output. Tulchinskii et al. (2024) proposes a persistent homology dimension estimator (PHD) that measures the intrinsic dimensionality of the manifold underlying the set of embeddings for a given text sample for deepfake text detection. Kushnareva et al. (2024) uses PHD to perform boundary detection (i.e., where the human author stops and the AI author begins) in the context of deepfake text detection. Lavery et al. (2024) performs fake news detection with a topologically-augmented BERT model. Proskura and Zaytsev (2024a) selects the best models to create an ensemble model with by first checking the topological similarity of the embeddings of the models. Das et al. (2021) extracts topological features using the sentence embeddings of BERT to represent texts as persistent image vectors to perform TEDtalk public speaking ratings classification. Gourgoulias et al. (2024) proposes an unsupervised approach for estimating the class separability of text datasets for binary and multi-class tasks. Rathore et al. (2023) proposes TopoBERT, a visual tool that topologically analyzes the fine-tuning process of the Transformer-based model using Mapper. Similarly, Byrne et al. (2022) uses Mapper for unsupervised topic modeling by representing the embeddings as a graph.
@@
Hidden Weights. Uchendu et al. (2024) proposes TopFormer, an end-to-end fine-tuned Topological RoBERTa-base model for deepfake text attribution by using the reshaped $pooled\_output$ (to 2D) as input for the TDA layer. Similar to Jakubowski et al. (2020), Triki (2021), and Shehu (2024), Garcia (2022) studies polysemous words. Using Mapper to visualize the hidden representations of polysemous words in the BERT transformer model, Garcia (2022) studies how BERT represents the semantics of ambiguous words. Similarly, Gardinazzi et al. (2024) proposes a novel metric - persistence similarity to describe the topological properties of internal representations in LLMs, detect topologically similar layers in an LLM, and prune redundant layers. Chauhan and Kaul (2022) proposes a novel scoring metric - persistence scoring function which captures the homology of the hidden representations of BERT, correlates with test accuracy, and is robust to adversarial perturbations. García-Castellanos et al. (2024) performs zero-shot model stitching by employing topological densification to encourage clustering of similar points. In addition, Ruppik et al. (2024) extracts topological features using persistent homology from RoBERTa’s hidden weights for dialogue term extraction. Finally, Kudriashov et al. (2024) probes BERT’s hidden weights using persistent homology when training BERT on new grammatical features, known as polypersonality.
@@
Attention Weights. Attention weights of transformers have also been utilized as input features to extract TDA features. The majority of the applications focus on two models: BERT and RoBERTa. Regarding BERT model, its attention weights have been transformed to both directed and undirected graphs, on top of which different TDA features are extracted that are useful for deepfake text detection Kushnareva et al. (2021) or investigation of whether language models learn grammatical concepts central to human linguistic competence (i.e., grammatical acceptability judgment) in different languages including English, Italian, Swedish, Russian Cherniavskii et al. (2022); Proskurina et al. (2023) and speech recognition task Tulchinskii et al. (2023). Similar techniques have also been used on RoBERTa’s attention weights for similar and other important tasks in linguistics such as grammatical checks Perez and Reinauer (2022), dialog term extraction Vukovic et al. (2022), deepfake text detection Uchendu et al. (2024), and document coherence Jain et al. (2024).
@@
We also notice a few works utilize the attention weights for analyzing different properties of NLP models such as model uncertainty estimation Kostenok et al. (2023), robustness evaluation of TDA features Perez and Reinauer (2022) and out-of-distribution detection (OOD) Pollano et al. (2024), dynamic weighting for ensemble models Proskura and Zaytsev (2024b). Especially,  Pollano et al. (2024) find that models utilizing TDA features extracted from attention weights perform best on far out-of-domain samples compared to when utilizing TDA features extracted from CLS outputs Pollano et al. (2024). This demonstrates that the choice of numerical representation used to extract TDA features is crucial.
@@
ELMo Embeddings. ELMo embeddings (Embeddings from Language Models) are a type of word representation that captures both the meaning of words and their usage in context Peters et al. (2018). Similar to other embeddings, ELMo has also been leveraged to extract topological features. Tymochko et al. (2021) uses ELMo embeddings to extract topological features for detecting fraudulent papers by examining their titles and abstracts. To compare the strength of ELMo embeddings, they further extract topological features with other embeddings (Word2Vec, GloVe, FastText, and Frequency Time Series) to ascertain the best embeddings to extract strong topological features.
@@ 4.2.4 Symbolic Representations
PLC refers to rules and methods used to encode letters, which can be part of various communication systems, cryptography techniques, or linguistic analyses. Letter coding transforms letters or characters into different symbols, numbers, or other forms. PSSC is similar to PLC but for extracting topological features from speech sounds. One particular application of PLC and PSSC is the study of tongue twisters. For instance, Yurchuk and Gurnik (2023) uses the PLC to create word embeddings for Ukrainian tongue twisters and extract topological features from such embeddings with persistent homology. This is to distinguish tongue twister from a simple narrative sentence using SVM and decision tree classifiers. Similarly, Kovaliuk et al. (2024) use PSSC for classifying spoken Ukrainian tongue twisters.
@@ 5.1 Theoretical Approaches Connecting TDA Features with Linguistic Principles
There is a need for theoretical approaches that better tie TDA features to linguistic phenomena. For instance, Draganov and Skiena (2024) investigates the shape of words and their embeddings in Indo-European languages and finds similar conclusions as Port et al. (2018) and Port et al. (2022), who investigate the syntactic topological space of such languages. They found that TDA features represent historical facts, such that the languages clustered closely together are similar or one influenced the other. Finally, we observe only 11 theoretical TDA approaches in NLP, compared to the 70+ non-theoretical approaches. Thus, we need more theoretical TDA approaches as it is impractical to achieve the depth and understanding of performance from a topological perspective without more investigations.
@@ 5.3 Novel Applications of TDA
Once more theoretical approaches of TDA are investigated, and issues barring the application of TDA on interpretable NLP tasks are mitigated, then we hope that TDA can be applied to even more novel, diverse and important tasks. From Section 4.1, we can see that TDA has been applied to 7 categories on non-theoretical NLP tasks. While many of the tasks are interesting, especially the speech & music processing and health applications, there are still nuanced niche fields that could benefit from TDA. One glaring application is on multi-lingual tasks, just as Haim Meirom and Bobrowski (2022) and García-Castellanos et al. (2024). Due to the benefits of TDA which include performing robustly on heterogeneous, imbalanced, and noisy data, its application on multi-lingual tasks is necessary. Other applications include: semantic and syntactic structure, forensic authorship, embedding space, and multi-modal (e.g., language-vision model) analysis

### v2

@@ Abstract
The surge of data available on the internet has led to the adoption of various computational methods to analyze and extract valuable insights from this wealth of information. Among these, the field of Machine Learning (ML) has thrived by leveraging data to extract meaningful insights. However, ML techniques face notable challenges when dealing with real-world data, often due to issues of imbalance, noise, insufficient labeling, and high dimensionality. To address these limitations, some researchers advocate for the adoption of Topological Data Analysis (TDA), a statistical approach that discerningly captures the intrinsic shape of data despite noise. Despite its potential, TDA has not gained as much traction within the Natural Language Processing (NLP) domain compared to structurally distinct areas like computer vision. Nevertheless, a dedicated community of researchers has been exploring the application of TDA in NLP, yielding 87 papers we comprehensively survey in this paper. Our findings categorize these efforts into theoretical and non-theoretical approaches. Theoretical approaches aim to explain linguistic phenomena from a topological viewpoint, while non-theoretical approaches merge TDA with ML features, utilizing diverse numerical representation techniques. We conclude by exploring the challenges and unresolved questions that persist in this niche field. Resources and a list of papers on this topic can be found at: https://github.com/AdaUchendu/AwesomeTDA4NLP.  [^1] DISTRIBUTION STATEMENT A. Approved for public release. Distribution is unlimited. This material is based upon work supported by the Department of the Air Force under Air Force Contract No. FA8702-15-D-0001. Any opinions, findings, conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the Department of the Air Force. © 2024 Massachusetts Institute of Technology. Delivered to the U.S. Government with Unlimited Rights, as defined in DFARS Part 252.227-7013 or 7014 (Feb 2014). Notwithstanding any copyright notice, U.S. Government rights in this work are defined by DFARS 252.227-7013 or DFARS 252.227-7014 as detailed above. Use of this work other than as specifically authorized by the U.S. Government may violate any copyrights that exist in this work.
@@ 1 Introduction
TDA aims to answer the main question - what is the true shape of a data? We survey 87 papers that have attempted to find an answer through various approaches. We categorize all observed approaches that incorporate TDA in NLP into two categories, namely theoretical Karlgren et al. (2014); Port et al. (2018) and non-theoretical Zhu (2013); Doshi and Zadrozny (2018). Theoretical approaches involve using TDA to explain linguistic phenomena by probing the topological space, shape, and evolution of topics. On the other hand, Non-theoretical approaches mainly discuss how to effectively apply existing numerical representation techniques in NLP to extract novel topological features with TDA. We will first discuss the principles behind TDA and the two main techniques employed for TDA feature extraction: Persistent Homology and Mapper.
@@ 2.2 Mapper
The intrinsic nature of the Mapper algorithm makes it advantageous in preserving structure even with mapping from one dimension to another. Furthermore, the clustering techniques allow it to be used to explain model performance as the clusters, and colors have meaning that can be further explored. Finally, Mapper is more useful for exploratory data analysis, while PH is more useful for analyzing point clouds and examining the persistence of features.
@@ 4.1 Tasks, Learning Types, Modality, and TDA techniques
Recently TDA has been heavily applied to deepfake text detection Løvlie (2023); Tulchinskii et al. (2024); Kushnareva et al. (2024, 2021); Uchendu et al. (2024), linguistic/grammatical acceptability classification Cherniavskii et al. (2022); Proskurina et al. (2023); Jain et al. (2024); Perez and Reinauer (2022), word sense induction and disambiguation Rawson et al. (2022); Jakubowski et al. (2020); Temčinas (2018), and polysemy word classification Jakubowski et al. (2020); Triki (2021); Shehu (2024); Garcia (2022). All of these five subtasks are particularly interesting because the unique labels have very similar features, making them non-trivial to distinguish from each other, making TDA a promising toolset, resulting in comparable performance to or even outperforming baselines.
@@ 4.2.2 Pre-trained Non-contextual Embeddings
Word2Vec embeddings are a type of word representation that allows words with similar meanings to have similar vector representations Mikolov et al. (2013). Word2vec has been used as inputs to persistent homology to extract a hierarchical document structure, known as story trees, that allows tracing storylines Haghighatkhah et al. (2022), or with mapper to construct a graph of word2vec embeddings to create a topological search engine Cornell (2020). TDA has also been applied on Word2vec to solve innovative problems such as measuring the distance between the literary style of Spanish poets belonging to similar genres Paluzo Hidalgo et al. (2019), the detection of fraudulent papers Tymochko et al. (2021), detecting topological loops in logical statements Tymochko et al. (2020), disease prediction from epidemic curves Petri and Leitao (2020), comparing between embeddings of text and speech Yessenbayev and Kozhirbayev (2022, 2024). TDA features extracted from Word2vec have also been used for unsupervised learning tasks such as topic modeling Holmes (2020), word sense induction and disambiguation Rawson et al. (2022); Temčinas (2018), derivation of the correlation between sentence vectors and their semantics Sun and Nelson (2023) or distinguishing subsets of Wikipedia data Wright and Zheng (2020). Such features have also been compared with those extracted from GloVe or LLMs’ embeddings using persistent homology Feng et al. (2024).
@@
FastText embeddings build upon the Word2Vec approach by incorporating subword information, enhancing the representation of rare words, and allowing for embedding out-of-vocabulary words Bojanowski et al. (2017). Jakubowski et al. (2020), Triki (2021), and Shehu (2024) use persistent homology to measure the polysemy of a word. They find that topological polysemy is inconsistent across different word embeddings. Jakubowski et al. (2020) further uses a similar technique to perform word sense induction & disambiguation from SemEval-2010 task. Also, Temčinas (2018) performs unsupervised word sense induction & disambiguation. Tymochko et al. (2021) extracts topological features from FastText embeddings to detect fraudulent papers.
@@ 4.2.3 Pre-trained Contextual Embeddings
CLS Embedding Output. Tulchinskii et al. (2024) proposes a Persistent Homology Dimension (PHD) estimator that measures the intrinsic dimensionality of the manifold underlying the set of CLS embeddings for a given text sample for deepfake text detection. PHD has also been utilized to perform boundary detection (i.e., where the human author stops and the AI author begins) in the context of deepfake text detection Kushnareva et al. (2024). CLS embeddings have been utilized to extract TDA features for other supervised learning tasks such as fake news detection Lavery et al. (2024), TEDtalk public speaking ratings classification Das et al. (2021) or unsupervised learning tasks such as topic modeling Byrne et al. (2022) and estimating class separability of text datasets Gourgoulias et al. (2024). Extracted topological features from CLS embeddings have also helped selecting the best ensemble model Proskura and Zaytsev (2024a) and visualizing the training process of transformer-based models Rathore et al. (2023).
@@
Hidden Weights. Uchendu et al. (2024) propose TopFormer, an end-to-end fine-tuned Topological RoBERTa-base model for deepfake text attribution by using the reshaped $pooled\_output$ (to 2D) as input for the TDA layer. Similar to Jakubowski et al. (2020), Triki (2021), and Shehu (2024), Garcia (2022) studies polysemous words. Using Mapper to visualize the hidden representations of polysemous words in the BERT transformer model, Garcia (2022) studies how BERT represents the semantics of ambiguous words. Similarly, Gardinazzi et al. (2024) proposes a novel metric - persistence similarity to describe the topological properties of internal representations in LLMs, detect topologically similar layers in an LLM, and prune redundant layers. Sun and Nelson (2023) derives the correlation between sentence vectors and their semantics using persistent homology. Chauhan and Kaul (2022) proposes a novel scoring metric - persistence scoring function which captures the homology of the hidden representations of BERT, correlates with test accuracy, and is robust to adversarial perturbations. García-Castellanos et al. (2024) performs zero-shot model stitching by employing topological densification to encourage clustering of similar points. In addition, Ruppik et al. (2024) extracts topological features using persistent homology from RoBERTa’s hidden weights for dialogue term extraction. Finally, Kudriashov et al. (2024) probes BERT’s hidden weights using persistent homology when training BERT on new grammatical features, known as polypersonality.
@@
Attention Weights. BERT’s attention weights have been transformed to both directed and undirected graphs, on top of which different TDA features are extracted for deepfake text detection Kushnareva et al. (2021), investigation of whether language models learn grammatical concepts central to human linguistic competence (i.e., grammatical acceptability judgment) in different languages including English, Italian, Swedish, Russian Cherniavskii et al. (2022); Proskurina et al. (2023) and speech recognition task Tulchinskii et al. (2023). Similar techniques have also been used on RoBERTa’s attention weights for similar and other important tasks in linguistics such as grammatical checks Perez and Reinauer (2022), dialog term extraction Vukovic et al. (2022), deepfake text detection Uchendu et al. (2024), and document coherence Jain et al. (2024).
@@
TDA features extracted from attention weights have also been used for model uncertainty estimation Kostenok et al. (2023), robustness evaluation of TDA features Perez and Reinauer (2022), and out-of-distribution detection (OOD) Pollano et al. (2024), dynamic weighting for ensemble models Proskura and Zaytsev (2024b). Especially,  Pollano et al. (2024) find that models that utilize TDA features extracted from attention weights perform best on far out-of-domain samples compared to those that use TDA features extracted from CLS outputs Pollano et al. (2024).
@@
ELMo Embeddings. ELMo embeddings are a type of word representation that captures both the meaning of words and their usage in context Peters et al. (2018). Similar to other embeddings, ELMo has also been leveraged to extract topological features. Tymochko et al. (2021) uses ELMo embeddings to extract topological features for detecting fraudulent papers by examining their titles and abstracts. To compare the strength of ELMo embeddings, they further extract topological features with other embeddings (Word2Vec, GloVe, FastText, and Frequency Time Series) to ascertain the best embeddings to extract strong topological features.
@@ 4.2.4 Symbolic Representations
PLC refers to rules and methods used to encode letters that fuel various communication systems, cryptography techniques, or linguistic analyses. Letter coding transforms letters or characters into different symbols, numbers, or other forms. PSSC is similar to PLC but for extracting topological features from speech sounds. One particular application of PLC and PSSC is the study of tongue twisters, such as for Ukrainian Yurchuk and Gurnik (2023); Kovaliuk et al. (2024).
@@ 5 Open Problems and Future Directions
5.1 LLMs to Convert TDA Concepts to Codes
One of the significant challenges in applying TDA to NLP tasks is the steep learning curve associated with its mathematical foundations, which are often accessible only to expert audiences. Moreover, the theorists who develop and understand these advanced concepts do not always collaborate with computational scientists to translate them into executable code. To address this gap, Liu et al. (2024) propose leveraging ChatGPT to generate Python code for TDA concepts by training it on these mathematical foundations. Their findings suggest that ChatGPT can alleviate this bottleneck, particularly for complex TDA concepts like hypergraphs, digraphs, and persistent harmonic space, which have not been as heavily explored as the vietoris-rips complex Liu et al. (2024). Similarly, experts can develop specialized code generators, such as fine-tuning models like Code-Llama Roziere et al. (2023) on TDA concepts. Creating a dedicated LLM for TDA code generation could significantly lower the entry barrier, encouraging the NLP community to explore TDA applications more innovatively.
@@ 5.2 Theoretical Approaches Connecting TDA Features with Linguistic Principles
There is a need for theoretical approaches that better tie TDA features to linguistic phenomena. For instance, Draganov and Skiena (2024) investigate the shape of words and their embeddings in Indo-European languages and finds similar conclusions as Port et al. (2018) and Port et al. (2022), who investigate the syntactic topological space of such languages. They found that TDA features represent historical facts as in languages clustered closely together are similar or influenced by the other. Finally, we observe only 11 theoretical TDA works in NLP, compared to the 70+ non-theoretical ones. Thus, we need more theoretical TDA approaches as it is impractical to achieve the depth and understanding of performance from a topological perspective without more investigations.
@@ 5.4 Novel Applications of TDA
When we have more theoretical approaches of TDA and issues barring the application of TDA on interpretable NLP tasks are mitigated, we can hope that TDA can be applied to even more novel, diverse and important tasks. From Section 4.1, we can see that TDA has been applied to 7 categories on non-theoretical NLP tasks. While many of the tasks are interesting, especially the speech & music processing and health applications, there are still nuanced niche fields that could benefit from TDA. One glaring application is on multi-lingual tasks, just as Haim Meirom and Bobrowski (2022) and García-Castellanos et al. (2024). Due to the benefits of TDA which include performing robustly on heterogeneous, imbalanced, and noisy data, its application on multi-lingual tasks is necessary. Other applications include: semantic and syntactic structure, forensic authorship, embedding space, and multi-modal (e.g., language-vision model) analysis